A carpet is to be installed in a room of length 8.90 m and width 6.1 m. Find the area of the room retaining the prope 106 Your r (b) Express the speed of light to five significant figures
o-10 points SerCP11 1.5.P 022 Find the height or length of these natural wonders in kilometers, meters, and centimeters. (a) a cave system with a mapped length of 365 miles km cm (b) a waterfall that drops 1,168 ft km cm (e) a 21,000 ft tall mountain km (d) a canyon with a depth of 73,000 ft km Noed Help? Read 9. 10/10 points Previous Answers SerCP11 1.5.P 024 O Type here to search 白さ00
Answer #1

4.

c = 2.99792458 x 108 m/s

a)

in three significant figures

c = 3.00 x 108 m/s

b)

c = 2.9979 x 108 m/s

c)

c = 2.997925 x 108 m/s

5.

a)

365 miles = 365 miles (1.60934 km /1 mile) = 365 x 1.60934 km = 587.4091 km

365 miles = 365 miles (1609.34 m /1 mile) = 365 x 1609.34 m = 587409.1 km

365 miles = 365 miles (160934 cm /1 mile) = 365 x 160934 cm = 58740910 cm

b)

1168 ft = 1168 ft (0.0003048 km/1 ft) = 1168 x 0.0003048 km = 0.3560064 km

1168 ft = 1168 ft (0.3048 m/1 ft) = 1168 x 0.3048 km = 356.0064 m

1168 ft = 1168 ft (30.48 cm/1 ft) = 1168 x 30.48 km = 35600.64 cm

c)

21000 ft = 21000 ft (0.0003048 km/1 ft) = 21000 x 0.0003048 km = 6.4008 km

21000 ft = 21000 ft (0.3048 m/1 ft) = 21000 x 0.3048 km = 6400.8 m

21000 ft = 21000 ft (30.48 cm/1 ft) = 21000 x 30.48 km = 640080 cm

d)

73000 ft = 73000 ft (0.0003048 km/1 ft) = 73000 x 0.0003048 km = 22.2504 km

73000 ft = 73000 ft (0.3048 m/1 ft) = 73000 x 0.3048 km = 22250.4 m

73000 ft = 116821000 ft (30.48 cm/1 ft) = 73000 x 30.48 km = 2225040 cm